OpenStudy (anonymous):

Simplify 15C3 A. 182 B. 455 C. 2,730 D. 910 *Ok alright I SWEAR this is the LAST one*

OpenStudy (anonymous):

Do you know how to do choose operation?

OpenStudy (anonymous):

No.. sorry :(

OpenStudy (anonymous):

Do you remember how to do P operation?

OpenStudy (anonymous):

v.v no.. :(

OpenStudy (anonymous):

why don't you learn how to do it?

OpenStudy (anonymous):

I tried to show you before.

OpenStudy (anonymous):

\[ ^nC_k = \frac{^nP_k}{k!} \]

OpenStudy (anonymous):

\[ k! = 1\times 2\times 3\times \ldots \times k \]

OpenStudy (anonymous):

\[ ^nP_k = \frac{n!}{(n-k)!} = n\times (n-1) \times (n-2)\times \ldots \times (n-k+1) \]

OpenStudy (anonymous):

ok I'm gonna try to figure this out.

OpenStudy (anonymous):

I got 455.

OpenStudy (anonymous):

:O I got it right on my test. THANK YOU!! :)